OpenStudy (anonymous):

The length of a rectangular garden is 9m more than the width if the perimeter is 156m, find the dimensions of the garden. With solution pls.

OpenStudy (saifoo.khan):

2[x+(x+9)]=156

OpenStudy (saifoo.khan):

2(2x+9) = 156 2x+18 = 78 2x = 78- 18 2x = 60 x = 30 width = 30 Length = 39

OpenStudy (anonymous):

4x+18 = 156 4x = 138 x = 138/4 x =34.5 Width = 34.5m Length = 43.5m Proof: 34.5*2 + 43.5*2 = 156